OpenStudy (anonymous):

giving medal! Solve . 6(c+4)-4c=-18 a. -4 c. 4 b. 21 d. -21

OpenStudy (anonymous):

6(c+4)-4c=-18 Expand 6(c+4) = 6c + 24, combine with -4c.

OpenStudy (caozeyuan):

6c+6*4-4c=-18, so 2c+24=-18, c=-21

OpenStudy (anonymous):

Wow first time I saw 5 or 6 people typing at the same time

OpenStudy (anonymous):

robz8 0 distribute the 6 6c+24-4c = -18 simplify 2c +24 = -18 subtract 24 from both sides 2c = -42 divide by 2x on both sides c = -21

OpenStudy (anonymous):

6c+24-4c=-18 2c+24=-18 2c=-42 c=-21

OpenStudy (caozeyuan):

cuz it's such easy question and you're giving a medal!

OpenStudy (amistre64):

medals are taken away from a deleted response ... which should entice the respondants to conform to our code of conduct better.

OpenStudy (caozeyuan):

@amistre64 , I'm second-language english, so I didn't get what you are saying